# A Blue Start: A large-scale pairwise and higher-order social network dataset

### v1

@@ IV Technical Validation
IV.1 Starter pack data
The starter pack data comprises 1,671,681 unique user and feed IDs and 301,322 starter packs. These statistics corroborate the findings of Ref. [29] which assembled a dataset of 335,416 starter packs. The discrepancy in sizes between these two datasets are attributable to the time of collection, the method of collecting the dataset, and the deletion of existing starter packs. Ref. [29] collected all of the starter packs from their introduction up to, and not including, January 1st, 2025, whereas with our method, we collected the list of DIDs on January 22nd, 2025 and queried the list of starter packs over several days ending on February 3rd, 2025. Out method relied on querying the static Bluesky API, whereas Ref. [29] used the Bluesky Firehose API and the Bluesky sync API (For more details, see Sec. II.5) to collect all starter packs. The advantage of the approach in Ref. [29] is that it can collect starter packs even if they are later deleted; it is estimated that 20% of all starter packs were deleted by January 1st, 2025 [29].
@@ IV.1 Starter pack data
We begin our analysis by examining the distribution of the number of starter packs that users belong to and create, as seen in Fig. 1(a). The starter pack sizes range from a minimum size of 0 to a maximum size of 4,070, with a mean of 32.86, a median of 24.0, and a mode of 8, while the number of starter packs to which a user account or feed belongs varies from 1 to 145,343 (with the highest count belonging to @bsky.app), averaging 5.92 memberships with a median of 1.0 and a mode of 1. Although the number of starter packs that users create varies widely, it is also quite homogeneous; while the majority (84.2%) of accounts in the network did not create a starter pack, those who did created 1.14 starter packs on average. Fig. 1 shows the hyperedge (starter pack) size distribution in Fig. 1(b) which, interestingly, is trimodal, with peaks at 8, 50, and 150. The peaks at 8 and 150 correspond to the minimum and maximum allowed starterpack sizes, and the peak at 50 is due to the Bluesky starter pack feature called ”Make one for me,” which pre-populates a starter pack with up to 50 accounts and feeds that a user follows. The distribution of connected component sizes (Figure 1, right) shows a very large giant connected component of 1.67M users accompanied by 382 components all smaller than the starter pack size limit of 150, the vast majority (361) of which are isolated starter packs.
Figure 2: Temporal starter pack statistics. Panel (a) plots the number of starter packs created each day and panel (b) plots the distribution of time elapsed between when a Bluesky account was created and when that account created a starter pack. The dashed lines in panel (a) illustrate the following notable events: (1) Bluesky announces the release of starter packs as a feature on June 26th, 2024 [62], (2) the Brazilian government bans the X platform on August 30th, 2024, (3) changes to the blocking feature on X are announced on October 16th, 2024 allowing blocked accounts to view posts of the blocking account, and (4) X announces new terms of service on November 15th, 2024 requiring users to consent to their posts being used to train X’s artificial intelligence models.
@@
Analyzing the times at which starter packs were created, we see four sharp peaks characterized by discontinuous jumps in the number of starter packs created and exponential decay after peak activity as illustrated in Fig. 2(a). These four peaks correspond to notable events in the history of both Bluesky and X. Much of Bluesky’s growth was spurred by user migration from X in response to platform and access changes. For example, in Fig. 2(a), we see (1) when starter packs were introduced as a feature [62], (2) when X was banned in Brazil, spurring an influx of Brazilian users, (3) the change to the blocking feature on X allowing blocked users to view the content of the account that blocked them, and (4) the change to the X terms of service, requiring users to consent to their posts being used to train X’s artificial intelligence models. Fig. 2(b) plots the distribution of time between when a user makes a Bluesky account and when they create a starter pack. This plot exhibits two interesting features: first, several starter packs are created by accounts with negative ages and second, the distribution of account ages is bimodal. Negative account ages are physically impossible; starter packs cannot be created prior to the creation of the associated user account. However, users have the ability to change their user data arbitrarily through the Bluesky API and this can cause data artifacts such as these. The distribution of account ages has a strong peak at zero and is bimodal, indicating that most users create a starter pack on the same day that they create an account and that early adopters of Bluesky are also active in creating starter packs.
Figure 3: $s$-line graph statistics of the starter pack network. The number of nodes and edges in the $s$-line graph of the starter pack network for all non-trivial values of $s$. In panels (a) and (b), the dashed line denotes the maximum starter pack size of 150.
@@
Line graph statistics are a powerful tool for analyzing higher-order network analysis and we use them to describe the edge overlap patterns in the starter pack network. A line graph of a higher-order network is a network with nodes corresponding to hyperedges and edges connecting two nodes if the two hyperedges that they represent share at least one node in common. Ref. [61] generalized this idea by introducing $s$-line graphs which only connect two nodes with an edge if the hyperedges they represent share at least $s$ nodes in common. We vary $s$, the edge overlap size, to get an idea of the strength of overlap across different regions of the network. This can be quite computationally intensive and it is often necessary to exploit algorithms such as those described in Ref. [63]. Because our analysis was confined to measuring the number of nodes and edges in the $s$-line graph, we were able to utilize a much more memory-efficient algorithm which focused on first finding cliques in the line graph and then looking for connections between edges in different cliques. This approach allowed us to avoid storing all the edges in a clique and simply store the node labels of the clique, which reduced the memory requirements considerably. Fig. 3 plots the number of nodes and edges in each $s$-line graph. Note that these graphs are extremely large and dense; for $s=1$, 29% of all possible edges are present in the line graph. Unsurprisingly, the maximum starter pack size is the most determinant of the connectedness of the line graph, followed by the size of the ”make one for me” starter pack feature corresponding to $s=50$. This higher-order network is far more dense and overlapping than many other existing higher-order datasets, providing an exciting opportunity for studies on higher-order network formation and higher-order algorithm development.
@@
Figure 4: Mesoscale starter pack statistics. Panel (a) plots the distribution of $k$-coreness, Panel (b) plots the probability of two user accounts appearing in $m$ different starter packs together, and Panel (c) plots the distribution of normalized entropy calculated for all starter packs. The illustrations of starter packs indicate at $s=0$, a starter pack completely contained in an inferred community and at $s=1$, a starter pack where every member belongs to a different community.
@@
We perform a hypergraph $k$-core decomposition [64] and, as seen in Fig. 4(a), the distribution has a very heavy tail. There are over 500 users with $k$-cores of over 1000 and this is indicative of many nodes deep in the core of the hypergraph, which implies the existence of large and intricate intersections between starter packs. Fig. 4(b) plots the co-occurrence distribution of user pairs, i.e., the probability of two user accounts appearing in $m$ different starter packs together. This can also be thought of as examining the distribution of weights in a weighted clique expansion of the hypergraph. From ocular regression, the distribution appears piecewise linear, suggesting three distinct regimes. There is a substantially high chance of having very large co-occurrence numbers. These insights suggest complex co-occurrence patterns, with a plethora of users participating in a multitude of starter packs simultaneously. In future studies, it will be fruitful to examine the co-occurrence of groups of $k$ nodes across different starter packs along the lines of techniques presented in [65]. To measure the community structure of the starter pack network, we run the Leiden [66] algorithm for two iterations on the unweighted pairwise projection of the largest connected component of the hypergraph to infer the community labels of all the nodes in the starter pack network, resulting in 654 clusters. The largest connected component was used so that isolated edges and small components would not inflate the number of clusters calculated. We use the inferred community labels, $\bm{\ell}_{e}=[\ell_{1},\dots,\ell_{k}]^{T}$, for each starter pack $e$ of size $k=|e|$, to specify an associated probability mass function $p_{e}(\lambda)=\sum_{1}^{k}\bm{1}_{\ell_{k}=\lambda}/k$. Then the normalized Shannon entropy for each each starter pack is defined as $s_{e}=-\sum_{\lambda}p_{e}(\lambda)\log_{2}[p_{e}(\lambda)]/\log_{2}(k)$, where $\lambda$. By this measure, the lowest entropy configuration ($s=0$) is a starter pack composed of nodes all belonging to the same community and the highest entropy configuration ($s=1$) is a starter pack composed of nodes all belonging to different communities. As seen in Fig. 4(c), there is a small spike at $s=0$ which can be traced to starter packs with 0, 1, or 2 members. Similarly, only starter packs with thirteen or fewer members achieved the maximum entropy configuration of $s=1$. This can be attributed to the constraints induced by a finite number of community labels; as the starter pack size $k$ increases, it becomes increasingly unlikely to find an edge composed of nodes, each with a different community label. However, Fig. 4(c) indicates that the vast majority of edges are closer to the maximum entropy configuration than to the minimum entropy configuration. This seems to indicate that the starter pack hypergraph exhibits relatively little community structure.
IV.2 Following data
@@
The following network consists of 26,692,837 distinct users and 1,654,721,704 directed following relationships linking a follower and a followed account. The minimum out-degree, or following count, in this network is 1 (a user who is not following anyone would not be included in the network), the maximum out-degree is 449,058, and the average out-degree is 64.75. The minimum in-degree, or follower count, in this network is 1, the maximum in-degree is 19,946,766, and the average in-degree is 92.47. The discrepancies in the mean degrees can be attributed to the fact that while 8,798,605 accounts have no followers, only 1,138,153 accounts do not follow any accounts. When including degree-zero nodes, the mean in- and out-degrees are both 61.99.
@@
Figure 5: Basic following network statistics. Panel (a) plots the in-degree and out-degree distributions and panel (b) plots the frequency of strongly and weakly connected component sizes.
@@
We begin our analysis by looking at the in- and out- degree distributions; both distributions are highly heterogeneous and heavy-tailed as seen in Fig. 5(a). In addition, we see that the in-degree distribution has a heavier tail than the out-degree distribution; not only are there more accounts following a large number of accounts, but the maximum number of accounts followed is larger than the number of followers of the most popular account. Fig. 5(b) plots the distributions of both the strongly and weakly connected components in the following network. The distribution of strongly connected component sizes contains one giant component of 16,594,500 nodes, along with several strongly connected components of 100 nodes or fewer. Looking at the distribution of weakly connected component sizes, we see an increase in the size of the giant component and a decrease in the number of small connected components, indicating that while the vast majority of the following graphs is strongly reachable, there are still ”in-components” and ”out-components” up to several hundred nodes in size. These components may or may not be reachable via snowball sampling [67] depending on the starting node of choice. A small number of components are truly disconnected from all other components via follower/following relationships.
IV.3 Comparing the starter pack and following network data
@@
The ”A Blue Start” dataset provides a bridge between pairwise and higher-order network representations. In Fig. 6, we compare the rankings induced by degree centrality in the starter pack hypergraph and the follows network. In the starter pack hypergraph, the degree centrality $d_{s}$ of a node (user) is the number of hyperedges (starter packs) of which the user is a member, while in the following network, the degree centrality $d_{f}$ of a node (user) is the sum of its following and followed by counts. Kendall’s $\tau_{\beta}$ rank correlation coefficient between two rankings ranges from -1 (total disagreement) to 1 (total agreement) and systematically handles ties in ranking. We start by computing $d_{s}$ and $d_{f}$ and only analyzing nodes appearing in both networks. We 1) rank the top $k$ nodes according to $d_{s}$ and compute the correlation between the ranks of these same nodes according to $d_{f}$, presenting the results as $d_{s}-d_{f}$; and 2) rank the top $k$ nodes according to $d_{f}$ and compute the correlation between the ranks of these same nodes according to $d_{s}$ and present the results as $d_{f}-d_{s}$. The fact that the maximal value of 1 is achieved at $k=1$ is not surprising, as the @bsky.app account is most central in both networks. However, the rankings quickly diverge, with a weaker correlation of approximately 0.5 for $k=10$ ($d_{s}-d_{f}$); this association increases to around 0.7 at $k=100$ and decreases with increasing $k$. Overall, relatively low correlation among the importance assigned to nodes in the higher-order (starter pack) and dyadic (following network) data demonstrates that the two contain complementary information about user importance, and that importance is conditioned on the context in which user interactions are measured.
@@ V Usage Notes
Processing the following network can be computationally intensive; The CSV file contains 1.3B lines of following relationships. Simply loading this dataset using the pandas Python package [68, 69] requires roughly 25 GB of memory. The size limitation of a pandas dataframe tends to be the amount of available memory and we encourage users of this dataset to verify that they have enough physical resources to process the following network data. To parallelize the processing of these data, we recommend loading the CSV with dask [70]. We opted to convert the dataset into a nested Python dictionary when computing the component sizes. The degree distributions were computed using a groupby operation in dask, and the component sizes were computed using the py-tarjan Python package.
@@
The starter pack data is a bit more manageable, requiring an order of magnitude less memory. However, as mentioned in Section IV.1, the line graph computed from the starter pack data is extremely dense; 29% of all possible pairs of starter packs share at least one member in common. When working with the edge connection data, one can dramatically decrease the number of shared starter pack members by removing the highest degree accounts. For example, the @bsky.app account is involved in 48% of all starter packs, and the top-5 most included accounts are all in greater than 10% of starter packs. In addition, one can decrease the number of edges by filtering out edges by their size, similar to the approach suggested in Ref. [40].
VI Data availability
All data is available on the Social Media Archive (SOMAR), which is a project by the Inter-university Consortium for Political and Social Research. The datasets described in this paper are available in the data collection titled ”A Blue Start: A large-scale pairwise and higher-order social network dataset” [71]. The dataset is available under SOMAR’s Controlled Download access method: researchers will be required to fill out a brief application detailing their research topic and agreeing to the ICPSR’s Terms of Use. This step is required in order to make sure the data are only used for research purposes; researchers must agree that they will not attempt to re-identify anonymized individuals.
@@ VI Data availability
The code used to analyze the starter pack and following networks is available on GitHub and at Ref. [72].

### v2

@@ Methods
Key Events
As we studied the temporal elements of the datasets (i.e. node creation, starter pack creation, following tie formation), we noticed spikes in activity that appear to be driven by key events; for example, a spike in node creation, starter pack creation, and following tie formation occurred when X/Twitter was banned in Brazil on August 30, 2024. We therefore include a table of these events; each event is indicated in the temporal plots, where relevant, using an integer (indicating an event that was important to multiple datasets) or a single letter (indicating an event pertinent only to the starter pack dataset) to identify it. Table 2 is color-coded to distinguish on-platform events from off-platform ones.
Identifier Event Date
1 Beta version of Bluesky opens; invitation codes are needed to join. February 17, 2023
2 Bluesky is open to the general public February 6, 2024
3 X blocks popular accounts in Brazil April 6, 2024
4 X formally allows adult and graphic content June 2, 2024
A Bluesky launches starter pack feature June 26, 2024
5 X is banned in Brazil August 30, 2024
6 New X terms of service allow blocked users to see posts October 16, 2024
7 New terms of service allow X to train AI on user content November 15, 2024
8 Donald Trump sworn in as U.S. president January 20, 2025
9 Nationwide Hands Off protests in U.S. April 5, 2025
Table 2: Significant events in the history of the Bluesky social media platform.
Nodes Count
Total users 39,650,447
Extant users 37,479,031
Users from listRepos 38,858,797
Active users 36,687,381
Inactive users 2,171,416
Taken down (moderation) 1,808,660
Deactivated 362,756
Users with no status 791,650
Table 3: User statistics.
Nodes
As seen in Fig. 2 and Table 3, the node dataset consists of 39.7 million accounts, of which 36.7 million are active, 2.17 million inactive, and 791k lacking an obtainable active/inactive status. Of the inactive accounts, 1.81 million had been taken down through moderation, and 363k had been deactivated.
Figure 2: Bluesky account creation and activity. Panel (a) plots the volume of account creations per day, overlaid with events significant to Bluesky’s history. Panel (b) plots the activity statuses for all accounts, with inactive accounts broken out into accounts taken down through moderation and deactivation.
Starter pack data
Starter Pack Network Statistic Value
Nodes 2,003,536
Edges 365,842
Components
Number 409
Largest 1,997,488 (99.7 of starter pack nodes)
Starterpacks created
Min / Max 1 / 252
Mean / Median 1.162 / 1
Mode 1
Users with none 84.3%
Starterpack Size
Min / Max 0 / 4,069
Mean / Median 34.724 / 24
Mode 50
Degree
Min / Max 1 / 175,159
Mean / Median 6.341 / 1
Mode 1
Top-5 account frequencies 47.88, 13.95, 11.59, 10.9, 10.47%
Line Graph
Density for $s=1\dots 5$ 0.2923, 0.0665, 0.034, 0.0207, 0.0139
k-core
k-core ($k\geq 1000$) size 772
Communities
Number 503
Top-5 community sizes 434,207; 333,148; 156,110; 126,567; and 98,434
Mean normalized edge entropy 0.16
Table 4: Starter pack network statistics
The starter pack data comprises 2,003,536 unique user and feed IDs and 365,842 starter packs. These statistics corroborate the findings of Ref. [5] which assembled a dataset of 335,416 starter packs. The discrepancy in sizes between these two datasets are attributable to the time of collection, the method of collecting the dataset, and the deletion of existing starter packs. Ref. [5] collected all of the starter packs from their introduction up to, and not including, January 1st, 2025, whereas with our method, we collected the list of DIDs on October 18, 2025 and queried the list of starter packs over several days ending on November 2, 2025. Out method relied on querying the Bluesky Sync API, whereas Ref. [5] used the Bluesky Firehose API and the Bluesky Sync API (For more details, see the section detailing limitations of our methodology) to collect all starter packs. The advantage of the approach in Ref. [5] is that it can collect starter packs even if they are later deleted; it is estimated that 20% of all starter packs were deleted by January 1st, 2025 [5].
@@ 7 New terms of service allow X to train AI on user content November 15, 2024
We begin our analysis by examining the distribution of the number of starter packs that users belong to and create, as seen in Fig. 3(a) and Table 4. Although the number of starter packs that users create varies widely, it is also quite homogeneous; while the majority (84.3%) of accounts in the network did not create a starter pack, those who did created 1.16 starter packs on average. Fig. 3 shows the hyperedge (starter pack) size distribution in Fig. 3(b) which, interestingly, is trimodal, with peaks at 8, 50, and 150. The peaks at 8 and 150 correspond to the minimum and maximum allowed starterpack sizes, and the peak at 50 is due to the Bluesky starter pack feature called “Make one for me,” which pre-populates a starter pack with up to 50 accounts and feeds that a user follows. The distribution of connected component sizes in Fig. 3(c) shows a giant connected component of 1.997M users accompanied by 408 components with all but two smaller than 1,000 users, the vast majority (383) of which are isolated starter packs.
Figure 4: Temporal starter pack statistics. Panel (a) plots the number of starter packs created each day and panel (b) plots the distribution of time elapsed between when a Bluesky account was created and when that account created a starter pack. The vertical lines in panel (a) correspond to relevant events enumerated in Table 2
@@
Analyzing the times at which starter packs were created, we see six sharp peaks characterized by discontinuous jumps in the number of starter packs created and exponential decay after peak activity as illustrated in Fig. 4(a). These peaks correspond to notable events in the history of Bluesky and X, as well as relevant U.S. political developments, which are detailed in Table 2. Fig. 4(b) plots the distribution of time between when a user makes a Bluesky account and when they create a starter pack. This plot exhibits two interesting features: first, a few starter packs were created by accounts with negative ages, and second, most starter packs are made by very recently created accounts. Negative account ages are physically impossible; starter packs cannot be created prior to the creation of the associated user account. However, users have the ability to change their user data arbitrarily through the Bluesky API and this can cause data artifacts such as these. We emphasize that the number of starter packs with erroneous time stamps is small; there are 8 starter packs with creation times prior to the creator’s account creation time, or 0.002% of all starter packs.
@@
Line graph statistics are a powerful tool for analyzing higher-order network analysis and we use them to describe the edge overlap patterns in the starter pack network. A line graph of a higher-order network is a network with nodes corresponding to hyperedges and edges connecting two nodes if the two hyperedges that they represent share at least one node in common. Ref. [1] generalized this idea by introducing $s$-line graphs which only connect two nodes with an edge if the hyperedges they represent share at least $s$ nodes in common. We vary $s$, the edge overlap size, to get an idea of the strength of overlap across different regions of the network. This can be quite computationally intensive and it is often necessary to exploit algorithms such as those described in Ref. [41]. Because our analysis was confined to measuring the number of edges in the $s$-line graph, we were able to utilize a much more memory-efficient algorithm which focused on first finding cliques in the line graph and then looking for connections between edges in different cliques. This approach allowed us to avoid storing all the edges in a clique and simply store the node labels of the clique, which reduced the memory requirements considerably. Fig. 5(a) plots the number of edges in each $s$-line graph. Note that these graphs are extremely large and dense; for $s=1$, 29% of all possible edges are present in the line graph. This is because several accounts are included in many starterpacks. For example, the @bsky.app account is involved in 48% of all starter packs, and the top-5 most included accounts are all in greater than 10% of starter packs. This higher-order network is far more dense and overlapping than many other existing higher-order datasets, providing an exciting opportunity for studies on higher-order network formation and higher-order algorithm development.
@@
Figure 5: Mesoscale starter pack statistics. Panel (a) plots the number of edges in the $s$-line graph of the starter pack network for all non-trivial values of $s$ (the dashed line denotes the maximum starter pack size of 150), Panel (b) plots the distribution of $k$-coreness, Panel (c) plots the probability of two user accounts appearing in $m$ different starter packs together, and Panel (d) plots the distribution of normalized entropy calculated for all starter packs. The illustrations of starter packs indicate at $s=0$, a starter pack completely contained in an inferred community and at $s=1$, a starter pack where every member belongs to a different community.
@@
We perform a hypergraph $k$-core decomposition [2] and, as seen in Fig. 5(b), the distribution has a very heavy tail. There are over 700 users with $k$-cores of over 1000 which is indicative of many nodes deep in the core of the hypergraph, implying the existence of large and intricate intersections between starter packs. Fig. 5(c) plots the co-occurrence distribution of user pairs, i.e., the probability of two user accounts appearing in $m$ different starter packs together. This can also be thought of as examining the distribution of weights in a weighted clique expansion of the hypergraph. From ocular regression, the distribution appears piecewise linear, suggesting three distinct regimes. There is a substantially high chance of having very large co-occurrence numbers. These insights suggest complex co-occurrence patterns, with a plethora of users participating in a multitude of starter packs simultaneously. In future studies, it will be fruitful to examine the co-occurrence of groups of $k$ nodes across different starter packs along the lines of techniques presented in Ref. [72].
To measure the community structure of the starter pack network, we run the Leiden [65] algorithm for two iterations on the unweighted pairwise projection of the largest connected component of the hypergraph to infer the community labels of all the nodes in the starter pack network, resulting in 503 clusters. The largest connected component was used so that isolated edges and small components would not inflate the number of clusters calculated. We use the inferred community labels to measure the Shannon entropy of each edge, normalized by the highest entropy configuration, where each node is in a different community. By this measure, the lowest entropy configuration ($s_{e}=0$) corresponds to a starter pack comprising nodes all belonging to the same community and the highest entropy configuration ($s_{e}=1$) corresponds to a starter pack comprising nodes all belonging to different communities. Fig. 5(d) indicates that the starter pack network contains strong community structure when compared to the distribution of normalized edge entropy values computed for a configuration model, with means of 0.16 and 0.576 respectively. These values are evidence of the correctness of the data; it is highly likely that the entropy values are much lower for an empirical dataset than for a random null model. This analysis also indicates that the network is likely suitable as a benchmark for community detection algorithms. In addition, even stronger community structure may also emerge if the network is filtered [34] to only include starter packs under a certain size or the network composed of starter packs with the most popular accounts removed.
Following data
Following Network Statistic Value
Nodes in network 36,447,725
Links in network 2,416,311,437
Largest connected component size
Strongly connected 20,495,220 (56.2% of following nodes)
Weakly connected 36,433,172 (99.96% of following nodes)
In-degree
Min / Max 1 / 28,062,787
Mean 105.8
Mean over all following nodes 66.3
Following nodes with in-degree 0 13,608,819
Out-degree
Min / Max 1 / 844,408
Mean 70.38
Mean over all following nodes 66.3
Followed nodes with out-degree 0 2,113,664
Table 5: Following network statistics
@@
As seen in Table 5, the following network consists of 36.4 million distinct users and 2.42 billion directed following relationships linking a follower and a followed account. The average out-degree for nodes following at least one account is 70.38, and the average in-degree for nodes that have at least one follower is 105.80. The discrepancies in the mean degrees can be attributed to the fact that while 13,608,819 accounts in the following network have no followers, only 2,113,664 accounts in the following network do not follow any accounts. When including degree-zero nodes, the mean in- and out-degrees are both 66.3.
@@
Figure 6: Basic following network statistics. Panel (a) plots the volume of following events for each day in our dataset, panel (b) plots the in-degree and out-degree distributions, and panel (c) plots the frequency of strongly and weakly connected component sizes. The vertical lines in panel (a) correspond to dates listed in Table 2
When plotting the volume of new following relationships over time in Fig. 6(a), we have omitted data with erroneous time stamps by filtering out following events with time stamps from before Bluesky was established as a platform. There are very few of these errors, however; 1 total follow has a verifiably incorrect time stamp. We also check for impossible timestamps where the date on which the follow event occurred is strictly less than the date on which either the follower or the followed account was created. Again, these errors are rare; 147,655 follow events fit this criterion, or 0.006% of follow events. As we discussed in the previous subsection, these impossible timestamps are likely a result of the fact that users can arbitrarily change their user data through the Bluesky API. The time series of number of follow events is characterized by exponential growth in the first six months followed by more consistent daily follow volume with intermittent spikes due to notable world and platform events. This approximate exponential growth can be attributed to the invite-only beta version of the app which allowed active users one new user per week with an invite code. As the daily volume of the new follows slowed down, several notable events characterized large changes in follow volume. For example, as illustrated in Fig. 6(a), (2) the public release of Bluesky on February 6, 2024 [53] and (3) the announcement of the blocking of certain popular accounts from Brazil [69] on April 5, 2024 sparked a large increase in new follows on the Bluesky platform. Similar to Fig. 4(a), we see the effect of several external events on the following behavior of Bluesky users, including changes in X’s terms of service and political events in the United States.
@@
We continue our analysis by looking at the in- and out- degree distributions; both distributions are highly heterogeneous and heavy-tailed as seen in Fig. 6(b). In addition, we see that the in-degree distribution has a heavier tail than the out-degree distribution; not only are there more accounts following a large number of accounts, but the maximum number of accounts followed is larger than the number of followers of the most popular account. Fig. 6(c) plots the distributions of both the strongly and weakly connected components in the following network. The distribution of strongly connected component sizes contains one giant component of 20,495,220 nodes (or 56.2% of the network), along with several strongly connected components of 100 nodes or fewer. Looking at the distribution of weakly connected component sizes, we see an increase in the size of the giant component to 36,443,172 users (or 99.96% of the network) and a decrease in the number of small connected components, indicating that while the vast majority of the following graphs is strongly reachable, there are still “in-components” and “out-components” up to several hundred nodes in size. These components may or may not be reachable via snowball sampling [26] depending on the starting node of choice. A small number of components are truly disconnected from all other components via follower/following relationships.
Historically, the created-at field returned by the static Bluesky API corresponded to the time at which the followed account was created. To verify that the time stamps of all following events truly correspond to following events, we calculate the standard deviation of the following times. A standard deviation of 0 for the in- and out-degree corresponds to follows all occurring on the same day for followed and following accounts, respectively.
Figure 7: Distribution of standard deviations of follow dates.
In Fig. 7, we see that the standard deviations are not only non-zero, but they differ as well. The distributions of these times make physical sense as well; it is more likely that a user follows most accounts within a small time range than for an account to be followed by many users within a small time range.
Comparing the starter pack and following network data
@@
The “A Blue Start” dataset provides a bridge between pairwise and higher-order network representations. In Fig. 8, we compare the rankings induced by degree centrality in the starter pack hypergraph and the follows network. In the starter pack hypergraph, the degree centrality $d_{s}$ of a node (user) is the number of hyperedges (starter packs) of which the user is a member, while in the following network, the degree centrality $d_{f}$ of a node (user) is the sum of its following and followed by counts. Kendall’s $\tau_{\beta}$ rank correlation coefficient between two rankings ranges from -1 (total disagreement) to 1 (total agreement) and systematically handles ties in ranking. We start by computing $d_{s}$ and $d_{f}$ and only analyzing nodes appearing in both networks. We 1) rank the top $k$ nodes according to $d_{s}$ and compute the correlation between the ranks of these same nodes according to $d_{f}$, presenting the results as $d_{s}-d_{f}$; and 2) rank the top $k$ nodes according to $d_{f}$ and compute the correlation between the ranks of these same nodes according to $d_{s}$ and present the results as $d_{f}-d_{s}$. The fact that the maximal value of 1 is achieved at $k=1$ is not surprising, as the @bsky.app account is most central in both networks. However, the rankings quickly diverge, with a weaker correlation of approximately 0.5 for $k=10$ ($d_{s}-d_{f}$); this association increases to around 0.7 at $k=100$ and decreases with increasing $k$. Overall, relatively low correlation among the importance assigned to nodes in the higher-order (starter pack) and dyadic (following network) data demonstrates that the two contain complementary information about user importance, and that importance is conditioned on the context in which user interactions are measured.
@@
Processing the following network can be computationally intensive; The CSV file contains 2.4B lines of following relationships and is 71 GB when inflated. Simply loading this dataset using the pandas Python package [46, 55] requires over 200 GB of memory. The size limitation of a pandas dataframe tends to be the amount of available memory and we encourage users of this dataset to verify that they have enough physical resources to process the following network data. To parallelize the processing of these data, we recommend loading the gzipped CSV or Parquet files with polars [56]. We opted to convert the dataset into a nested Python dictionary when computing the component sizes. The degree distributions were computed using a value_counts operation in polars, and the component sizes were computed using the py-tarjan Python package. Both of these computations required 256 GB of memory to run successfully.
Generally speaking, we recommend against using networkx to work with the following network; loading the following network with igraph or graph-tool consumes relatively tractable amounts of memory, but doing the same with networkx used over 500 GB of memory before even loading the full network. We provide example code to load the following network as a graph-tool or igraph object; for those who prefer to use dataframes to work with the data, we also include a copy of the following network in Parquet format. Note that we include functionality for converting follow dates to integer edge attributes (i.e. days since the epoch) as well as code to load the node and starter pack dataset with corresponding conversions of all relevant dates to integers. This minimizes the dataset’s memory footprint while still making it possible to work with the temporal aspects of the dataset.
When working with igraph and graph-tool, we find that loading the edges in large batches provides the best balance of minimizing memory usage and loading the graph in a timely fashion. If the entire graph is loaded as a polars dataframe and then converted into a graph-tool or igraph object, RAM usage also will quickly exceed 500 GB. For this reason, we recommend loading around 5-10% of the edges at a time, streaming data from the csv.gz file, then inserting them into the graph object. Parsing and inserting 100 million edges consistently took around 6 minutes for the graph-tool loading process. For igraph, adding edges takes longer as the graph grows. Parsing and inserting the first 250 million edges took around 15 minutes, while the last set of edges, numbering less than 250 million, required 35 minutes. Overall, loading the graph as an igraph object used 460 GB of RAM and took 5.5 hours on an HPC instance, while loading the graph into graph-tool consumed 310 GB of RAM and took 2.75 hours.
@@
The starter pack data is a bit more manageable, requiring an order of magnitude less memory. However, as mentioned in the section on data validation, the line graph computed from the starter pack data is extremely dense; 29% of all possible pairs of starter packs share at least one member in common. When working with the edge connection data, one can dramatically decrease the number of shared starter pack members by removing the highest degree accounts.
As noted when discussing the validity of the data, a small percentage of the starter packs and following relationships have incorrect time stamps. In the case of the starter packs, we recommend either discarding the packs with incorrect dates, setting the times to the creator’s account creation date, or sampling an account age at the time of starter pack creation by bootstrapping the distribution of realizable account ages presented in Fig. 4(b). In the case of the following relationships, we recommend either discarding the follow events that have impossible dates or setting the follow dates to the maximum of the follower and followed accounts’ creation dates.
Data Availability
The dataset is hosted on the Social Media Archive (SOMAR) at the Inter-university Consortium for Political and Social Research (ICPSR) at https://doi.org/10.3886/ICPSR300499.
Code Availability
@@
The code used to analyze the starter pack and following networks is available on GitHub
(https://github.com/nwlandry/a-blue-start) and at Ref. [61].